In a triangle, the measure of one angle is 60 degrees, the second angle is twice the measure of the third angle. Find the measures of the second and third angles.
Answer by checkley75(3666)

60+2*X+X=90
60+2X+X=90
3X=90-60
3X=30
X=30/3
X=10 DEGREES FOR THE THIRD ANGLE.
2*10=THE SECOND ANGLE.
PROOF
60+20+10=90
90=90